# Spurious Correlations in Machine Learning: A Survey

### v1

@@ 3.3 Learning Strategy
3.3.1 Optimization-based Methods
In tackling spurious correlations, a range of optimization-based methods have been developed. Optimization methods focus on proposing reasonable training objectives for the models to learn more reasonable metrics, thus improving the model’s robustness and fairness. Group Distributionally Robust Optimization (Group DRO) Sagawa et al. (2019) is the very first work, which focuses on minimizing the worst-case loss across different predefined groups in a dataset. This method is effective in ensuring that the model performs well across various subpopulations, reducing the risk of overfitting to dominant trends. Conditional Value at Risk (CVaR) DRO Levy et al. (2020) extends this concept by optimizing the expected loss over the worst-case distribution of the data. This approach is useful for managing risk in scenarios where extreme cases can have significant impacts, thereby promoting a more robust model performance.
Instead of following the worst-case loss from Group DRO, Liu et al. (2023) proposes a new logit correction (LC) loss by adjusting the model’s output logits based on group membership, aiming to balance performance across different groups. This correction is particularly useful in scenarios where certain groups are underrepresented or more prone to misclassification due to biased data.
3.3.2 Ensemble Learning
Ensemble learning denotes a set of methods that ensembles multiple independently trained biased models to make a more generalizable unbiased model. For instance, the LWBC method, as described by Kim et al. (2022), utilizes a committee of classifiers to identify bias-conflicting data—those free from spurious correlation—and emphasizes these during the training of the main classifier. Similarly, Bahng et al. (2020) introduced ReBias, a framework for training de-biased representations by making them statistically independent from a set of intentionally biased representations.
Several methods also consider a two-stage ensemble strategy. Lee et al. (2023) proposed DivDis, a two-stage framework that initially generates a variety of hypotheses using unlabeled data and subsequently selects the most suitable one with minimal additional supervision. Additionally, the Learning from Failure (LfF) Nam et al. (2020) approach involves a two-step process where the first network is intentionally trained with bias, and the second network is debiased by focusing on samples that counteract this bias. Collectively, these methods exemplify the diverse strategies in ensemble learning aimed at mitigating spurious correlations in machine learning models.
3.3.3 Identification then Mitigation
@@
The following methods consider identifying the underperforming samples in the dataset and then mitigating the spurious correlations based on previously identified data. Just Train Twice (JTT) Liu et al. (2021) proposes a simple yet effective strategy. It initially identifies examples that are hard to classify, then re-trains the model to focus specifically on these challenging tasks by upweighting the identified points. First Amplify Correlations and Then Slice (FACTS) Yenamandra et al. (2023) initially amplifies correlations to fit a simplified bias-aligned hypothesis through empirical risk minimization. Then it applies correlation-aware slicing, using mixture modeling in the bias-aligned feature space to identify and address underperforming data slices representing distinct correlations. SePArate early and REsample (SPARE) Yang et al. (2023a) operates by distinguishing large groups with spurious correlations early in the training process. It then employs importance sampling to balance these group sizes, effectively alleviating the influence of spurious correlations.
Other methods aim to quantify spurious correlations as scores in the dataset before taking corresponding mitigative actions. Less-Learn-Shortcut (LLS) Du et al. (2023) focuses on quantifying the degree of bias in samples and proportionately down-weighting them. This method allows for a more balanced learning process by reducing the influence of heavily biased data. Spuriosity Rankings Moayeri et al. (2023) involves sorting data within each class according to their level of spuriosity, as determined by neural features in an interpretable network. This technique facilitates the efficient retrieval of natural examples and counterfactuals related to relevant spurious cues from large datasets, thereby enabling effective measurement and mitigation of reliance on spurious features.
Additionally, there is also one method Yadav et al. (2022) which transforms the natural language data to a rule-based logic formulation and then uses logical negation to address spurious correlations. The proposed Tsetlin Machine (TM) employs a rule-based architecture that learns both simple and complex correlations by combining features and their negations. This approach enables the extraction of rule-based reasoning from TM, yielding explainable AND-rules that incorporate both negated and non-negated elements.
@@
3.3.4 Finetuning Strategy
Finetuning strategies focus on refining existing general prediction models by selectively adjusting either a portion of the model or the data used. This selective finetuning often targets outlier data and crucial interpretable parts of the model, aiming to address the challenge of spurious correlations. MaskTune Asgari et al. (2022) encourages a model to consider a broader range of input variables, mapping them to the same target, thereby promoting a more comprehensive understanding of the data. Influence Tuning Han and Tsvetkov (2021) takes a unique approach by backpropagating a target instance’s attribution interpretation directly into the model, allowing for more targeted adjustments. Explanation-Based Finetuning Ludan et al. (2023) involves training the model on artificially constructed datasets that include various types of spurious cues, and subsequently testing it on sets devoid of these cues.
Recently, some methods only consider a relatively simple strategy, last-layer finetuning, while they achieve surprisingly good performance. Deep Feature Reweighting (DFR) Kirichenko et al. (2023) focuses on retraining the last layer of a network to reduce the impact of spurious features, enhancing the model’s worst-group performance and reducing reliance on non-essential features, like background and texture. Selective Last-Layer Finetuning (SELF) LaBonte et al. (2023) employs a strategy of finetuning the last layer using a small, more group-balanced dataset, showing improved performance to near-DFR levels with no group annotations.
3.3.5 Adversarial Training
Adversarial training strengthens a model against adversarial inputs, enhancing its ability to generalize and learn on general data patterns, thereby reducing the effects of spurious correlations. Backdoor Poisoning Attacks He et al. (2023) is an adversarial attack method where the attacker injects several watermarked, mislabeled training examples into a training set. This paper argues that backdoor poisoning attacks exhibit spurious correlations between simple text features and classification labels, and accordingly proposes methods for mitigating spurious correlation as a means of defense.
@@ 3.4 Other Methods
This category collects the methods with unique problem settings. While these techniques may not be universally applicable, they provide insightful solutions for mitigating spurious correlations within their specialized fields. Test-Time Label-Shift Adaptation (TTLSA) Sun et al. (2023) addresses label shifts occurring at test time. MultiTask Causal Representation Learning (MT-CRL) Hu et al. (2022) enhances learning in multi-task environments by identifying causal relationships. RSC-MDP Ding et al. (2023) offers a novel perspective in reinforcement learning settings, focusing on reducing spurious dependencies for more reliable decision-making processes. Another work Yang et al. (2023b) also focuses on applying multi-modal models to leverage different modalities to explicitly set apart spurious attributes from the affected class.
@@ 4.1 Datasets
In this subsection, we briefly introduce the popular datasets for spurious correlation in machine learning (Table 1). Under the spurious correlation settings, the correlation between the label $y$ and the spurious attribute $a$ at training time does not hold attest time. Most of benchmarks are constructed from these existing datasets. Vision tasks are the most common used benchmarks to understand spurious correlations, such as Colored MNIST Lee et al. (2021), Corrupted CIFAR-10 Lee et al. (2021), CelebA Sagawa et al. (2019), Waterbirds Sagawa et al. (2019), subset of ImageNet Hendrycks and Dietterich (2018); Xiao et al. (2020); Hendrycks et al. (2021), FMOW Christie et al. (2018), BAR Nam et al. (2020), NICO He et al. (2021), MetaShift Liang and Zou (2021), Spawrious Lynch et al. (2023), and 2 VQA benchmarks Agarwal et al. (2020). In NLP, CivilComments Borkan et al. (2019), MultiNLI Sagawa et al. (2019), and FDCL18 Founta et al. (2018) are also widely used benchmarks. Apart from vision and language areas, spurious correlation is also an important research topic for AI in healthcare. For instance, ChestX-ray14 Wang et al. (2017), CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), MIMIC CXR Johnson et al. (2019) are widely used chest X-ray datasets to understand spurious correlations in medical imaging.
@@ 4.2 Metrics
The most common metric to evaluate the robustness of a model $f_{\theta}$ to spurious correlations is the worst-group-accuracy $\operatorname{Acc}_{\text{wg}}(f_{\theta})$ Idrissi et al. (2022); Chaudhuri et al. (2023), which is defined as the minimum accuracy across groups of test samples, i.e.,
@@ 5 Discussion
Refer to caption Hospital tags
Refer to caption Stripes
Refer to caption Medical devices
@@ 5.3 How will Foundation Models Affect this Area?
Another promising direction is to study the spurious correlations in foundation models. As foundation models are overparameterized models, a critical concern is that overparameterization in these models may exacerbate spurious correlations Sagawa et al. (2020). This suggests that while being powerful, foundation models may also enhance unintended or misleading spurious associations, especially on previously unseen data. The hallucination of Large Language Models (LLMs) is a well-acknowledged example, where LLMs occasionally generate spurious content that diverges from user inputs, contradicts previously generated context, or misaligns with established factual knowledge. New methods, such as introducing knowledge graphs into foundation models, may be developed to address this issue.

### v2

@@ 3.3 Learning Strategy
Optimization-based Methods
In tackling spurious correlations, a range of optimization-based methods have been developed. Optimization methods focus on proposing reasonable training objectives for the models to learn more reasonable metrics, thus improving the model’s robustness and fairness. Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is the very first work, which focuses on minimizing the worst-case loss across different predefined groups in a dataset. This method is effective in ensuring that the model performs well across various subpopulations, reducing the risk of overfitting to dominant trends. Conditional Value at Risk (CVaR) DRO (Levy et al., 2020) extends this concept by optimizing the expected loss over the worst-case distribution of the data. This approach is useful for managing risk in scenarios where extreme cases can have significant impacts, thereby promoting a more robust model performance.
Instead of following the worst-case loss from Group DRO, (Liu et al., 2023) proposes a new logit correction (LC) loss by adjusting the model’s output logits based on group membership, aiming to balance performance across different groups. This correction is particularly useful in scenarios where certain groups are underrepresented or more prone to misclassification due to biased data.
Ensemble Learning
Ensemble learning denotes a set of methods that ensembles multiple independently trained biased models to make a more generalizable unbiased model. For instance, the LWBC method, as described by (Kim et al., 2022), utilizes a committee of classifiers to identify bias-conflicting data—those free from spurious correlation—and emphasizes these during the training of the main classifier. Similarly, (Bahng et al., 2020) introduced ReBias, a framework for training de-biased representations by making them statistically independent from a set of intentionally biased representations.
Several methods also consider a two-stage ensemble strategy. (Lee et al., 2023) proposed DivDis, a two-stage framework that initially generates a variety of hypotheses using unlabeled data and subsequently selects the most suitable one with minimal additional supervision. Additionally, the Learning from Failure (LfF) (Nam et al., 2020) approach involves a two-step process where the first network is intentionally trained with bias, and the second network is debiased by focusing on samples that counteract this bias. Collectively, these methods exemplify the diverse strategies in ensemble learning aimed at mitigating spurious correlations in machine learning models.
Identification then Mitigation
@@
The following methods consider identifying the underperforming samples in the dataset and then mitigating the spurious correlations based on previously identified data. Just Train Twice (JTT) (Liu et al., 2021) proposes a simple yet effective strategy. It initially identifies examples that are hard to classify, then re-trains the model to focus specifically on these challenging tasks by upweighting the identified points. First Amplify Correlations and Then Slice (FACTS) (Yenamandra et al., 2023) initially amplifies correlations to fit a simplified bias-aligned hypothesis through empirical risk minimization. Then it applies correlation-aware slicing, using mixture modeling in the bias-aligned feature space to identify and address underperforming data slices representing distinct correlations. SePArate early and REsample (SPARE) (Yang et al., 2023a) operates by distinguishing large groups with spurious correlations early in the training process. It then employs importance sampling to balance these group sizes, effectively alleviating the influence of spurious correlations.
Other methods aim to quantify spurious correlations as scores in the dataset before taking corresponding mitigative actions. Less-Learn-Shortcut (LLS) (Du et al., 2023) focuses on quantifying the degree of bias in samples and proportionately down-weighting them. This method allows for a more balanced learning process by reducing the influence of heavily biased data. Spuriosity Rankings (Moayeri et al., 2023) involves sorting data within each class according to their level of spuriosity, as determined by neural features in an interpretable network. This technique facilitates the efficient retrieval of natural examples and counterfactuals related to relevant spurious cues from large datasets, thereby enabling effective measurement and mitigation of reliance on spurious features.
Additionally, there is also one method (Yadav et al., 2022) which transforms the natural language data to a rule-based logic formulation and then uses logical negation to address spurious correlations. The proposed Tsetlin Machine (TM) employs a rule-based architecture that learns both simple and complex correlations by combining features and their negations. This approach enables the extraction of rule-based reasoning from TM, yielding explainable AND-rules that incorporate both negated and non-negated elements.
@@
Finetuning Strategy
Finetuning strategies focus on refining existing general prediction models by selectively adjusting either a portion of the model or the data used. This selective finetuning often targets outlier data and crucial interpretable parts of the model, aiming to address the challenge of spurious correlations. MaskTune (Asgari et al., 2022) encourages a model to consider a broader range of input variables, mapping them to the same target, thereby promoting a more comprehensive understanding of the data. Influence Tuning (Han & Tsvetkov, 2021) takes a unique approach by backpropagating a target instance’s attribution interpretation directly into the model, allowing for more targeted adjustments. Explanation-Based Finetuning (Ludan et al., 2023) involves training the model on artificially constructed datasets that include various types of spurious cues, and subsequently testing it on sets devoid of these cues.
Recently, some methods only consider a simple strategy, last-layer finetuning, while they achieve good performance. Deep Feature Reweighting (DFR) (Kirichenko et al., 2023) focuses on retraining the last layer of a network to reduce the impact of spurious features, enhancing the model’s worst-group performance and reducing reliance on non-essential features, like background and texture. Selective Last-Layer Finetuning (SELF) (LaBonte et al., 2023) employs a strategy of finetuning the last layer using a small, more group-balanced dataset, showing improved performance to near-DFR levels with no group annotations.
Adversarial Training
Adversarial training strengthens a model against adversarial inputs, enhancing its ability to generalize and learn on general data patterns, thereby reducing the effects of spurious correlations. Backdoor Poisoning Attacks (He et al., 2023) is an adversarial attack method where the attacker injects several watermarked, mislabeled training examples into a training set. This paper argues that backdoor poisoning attacks exhibit spurious correlations between simple text features and classification labels, and accordingly proposes methods for mitigating spurious correlation as a means of defense.
@@ 3.4 Other Methods
This category collects the methods with unique problem settings. While these techniques may not be universally applicable, they provide insightful solutions for mitigating spurious correlations within their specialized fields. Test-Time Label-Shift Adaptation (TTLSA) (Sun et al., 2023) addresses label shifts occurring at test time. MultiTask Causal Representation Learning (MT-CRL) (Hu et al., 2022) enhances learning in multi-task environments by identifying causal relationships. RSC-MDP (Ding et al., 2023) offers a novel perspective in reinforcement learning settings, focusing on reducing spurious dependencies for more reliable decision-making processes. Another work (Yang et al., 2023b) also focuses on applying multi-modal models to leverage different modalities to explicitly set apart spurious attributes from the affected class.
@@ 4.1 Datasets
In this subsection, we briefly introduce the popular datasets for spurious correlation in machine learning (Table 1). Under the spurious correlation settings, the correlation between the label $y$ and the spurious attribute $a$ at training time does not hold attest time. Most of benchmarks are constructed from these existing datasets. Vision tasks are the most common used benchmarks to understand spurious correlations, such as Colored MNIST (Lee et al., 2021), Corrupted CIFAR-10 (Lee et al., 2021), CelebA (Sagawa et al., 2019), Waterbirds (Sagawa et al., 2019), subset of ImageNet (Hendrycks & Dietterich, 2018; Xiao et al., 2020; Hendrycks et al., 2021), FMOW (Christie et al., 2018), BAR (Nam et al., 2020), NICO (He et al., 2021), MetaShift (Liang & Zou, 2021), Spawrious (Lynch et al., 2023), SpuCo (Joshi et al., 2023), and 2 VQA benchmarks (Agarwal et al., 2020). In NLP, CivilComments (Borkan et al., 2019), MultiNLI (Sagawa et al., 2019), and FDCL18 (Founta et al., 2018) are also widely used benchmarks. Apart from vision and language areas, spurious correlation is also an important research topic for AI in healthcare. For instance, ChestX-ray14 (Wang et al., 2017), CheXpert (Irvin et al., 2019), PadChest (Bustos et al., 2020), MIMIC CXR (Johnson et al., 2019) are widely used chest X-ray datasets to understand spurious correlations in medical imaging.
@@ 4.2 Metrics
The most common metric to evaluate the robustness of a model $f_{\theta}$ to spurious correlations is the worst-group-accuracy $\operatorname{Acc}_{\text{wg}}(f_{\theta})$ (Idrissi et al., 2022; Chaudhuri et al., 2023), which is defined as the minimum accuracy across groups of test samples, i.e.,
@@ 5 Discussion
Hospital tags
Stripes
Medical devices
@@ 5.3 How will Foundation Models Affect this Area?
Another promising direction is to study the spurious correlations in foundation models. As foundation models are overparameterized models, a critical concern is that overparameterization in these models may exacerbate spurious correlations (Sagawa et al., 2020). This suggests that while being powerful, foundation models may also enhance unintended or misleading spurious associations, especially on previously unseen data. The hallucination of Large Language Models (LLMs) is a well-acknowledged example, where LLMs occasionally generate spurious content that diverges from user inputs, contradicts previously generated context, or misaligns with established factual knowledge. New methods, such as introducing knowledge graphs into foundation models, may be developed to address this issue.